OpenStudy (anonymous):

Differentiate: f(x)=(x^2+3x)e^(x^2)

OpenStudy (anonymous):

product rule would work nicely \[(fg)'=f'g+g'f\] with \[f(x)=x^2+3x,f'(x)=2x+3,g(x)=e^x, g'(x)=e^x\] plug and chug